OpenStudy (anonymous):

This problem was asked and closed before it is solved. I will ask it here again so it can hopefully be addressed Solve \[ \frac {dx}{y+z} =\frac {dy}{x+z}=\frac {dz}{x+y} \]

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (freckles):

hey @eliassaab do you have one of those computer answers you always give :) I would like to see the answer

ganeshie8 (ganeshie8):

*

OpenStudy (freckles):

Hey I found something that might be useful to look at. I'm pretty sure it is above what I can comprehend.

OpenStudy (freckles):

Like I was thinking a system would come in handy but I didn't know how to set it up... I also don't know how he got those equations he got. x'=y+z and so on...

OpenStudy (anonymous):

I will post the solution after some responses.

OpenStudy (anonymous):

@freckles, well for the solution above, the computer generated the pdf file from a latex file!!!

OpenStudy (anonymous):

On of the solutions found above is graphed in the attached as the the intersections of two surfaces.

OpenStudy (anonymous):

@ganeshie8

ganeshie8 (ganeshie8):

Cleverly done sir xD

OpenStudy (anonymous):

Complete solution including the elimination of t.

OpenStudy (fibonaccichick666):

wow, thanks guys! Turns out it was PDE. The teacher is just a jerk. I plan on using that solution to study though elias! Thank you everyone!

OpenStudy (anonymous):

YW @FibonacciChick666

OpenStudy (fibonaccichick666):

you want me to post the solution the teacher said?

OpenStudy (anonymous):

yes. I would love to see it.

OpenStudy (fibonaccichick666):

Note that \[\frac{dx}{z+y}=\frac{dy}{x+z}=\frac{dz}{x+y}=\frac{d(x-y)}{y-x}=\frac{d(x-z)}{z-x}=\frac{d(y-z)}{z-y}\] Hence we obtain solution via first integrals.\[ln|x-y|-ln|y-z|=C_1~~~~~~ln|x-z|-ln|y-z|=C_2\] This gives us two first integrals \[\Phi (x,y,z):=\frac{x-y}{y-z},~~~~~~~\Psi (x,y,z):=\frac{x-z}{y-z}.\] The general first integral is \[\Phi (x,y,z):=F(\frac{x-y}{y-z},\frac{x-z}{y-z}).\] Where F(u,v) is an arbitrary function. Indeed \(\Phi\) is the general solution of \[(z+y)\frac{\delta \Phi}{\delta x}+(x+z)\frac{\delta \Phi}{\delta y}+(x+y)\frac{\delta \Phi}{\delta z}=0.\] End

OpenStudy (fibonaccichick666):

He's a jerk. And that word is not strong enough

OpenStudy (anonymous):

No, he is not a jerk. Actually, I can deduce my solution from his method.

OpenStudy (fibonaccichick666):

To put something on a test we have not even learned that is not in the spectrum of the class? Yes he's a jerk.

OpenStudy (fibonaccichick666):

we never learned what a first integral even is

OpenStudy (fibonaccichick666):

Then tell us you can ask anyone for help on this, encouraged us to ask for help even, from other teachers, graduate students, anyone, so long as they are not our classmates or him-yes, he is a jerk. then to teach above our heads and say halfway through lecture "Oh, I forgot you are not graduate students." not to mention giving less than 48 hours to complete 37 problems of this caliber. Jerk worthy.